Tom and Mark are stranded on an island. If Tom rounds up horses all day, he can lasso

12 horses. If Tom catches pigs all day, the maximum he can catch is 24. He can also produce any linear combination of

the animals above. Mark can round up 36 if he rounds up horses all day. Mark can catch 18 pigs if he

catches pigs all day. Mark, like Tom, can produce any linear combination of horses and pigs.

1. Draw Tom's PPF.

2.Who is the low cost producer of pigs? How do you know?

3.Draw the PPF of the island's economy under autarky and with exchange. Be sure to label the important points.

4.Depict the marginal cost (or supply) of horses on the island

5.What is the range of potentially acceptable exchange rates for horse?

6.If they agree their combined demand for horses is 24, what is the maximum number of pigs they can catch in a day?
Answer #1

1. the island there are only two people Tom and Mark and they can not get away from there. Now if Tom rounds up horses for all day Tom can lasso 12 horses. If Tom cathches pigs for all day then maximum pigs that Tom can cathes is 24. So either Tom can lasso 12 horses in a whole day or can cathes 24 pigs in whole day. So Tom's PPF will be based on a vertical intercept and horizontal intercept. If we measure horses in vertical axis and pigs in horizontal axis. Then vertical intercept will be 12 and horizontal intercept will be 24.

12. ME

2. Tom can either lasso 12 horses or can pick 24 pigs in whole day. So Tom's relative caost for 1 pig is 12/24 horse =1/2 horse. Now Mark can lasso 36 horses or can pick 18 pigs in whole day. Now Mark's relative cost for 1 pig is 36/18 horses = 2 horses. By calculating relative cost for 1 pig in case of both Tom and Mark we can say relative cost is lower for Tom. Because Tom has to sacrifice 1/2 horse for 1 pig but Mark has to sacrifice 2 horses for 1 pig.

3. Under auterky situation PPF of Tom is such that it can produce maximum 12 horses or 24 pigs. Under auterky PPF of Mark is such that it can produce 36 horses or 18 pigs. So total in the economy both can together 36 horses or 24 pigs. But when they will involve in trade they can produce 36 horse and 24 pigs. It means they can produce either 48 horse (as 24 pigs =12 horse) or 42 pigs (as 36 horse = 24 pigs). So without exchange vertical intercept ( measuring horse in vertical axis) will be 36 and horizontal intercept will be 24( measuring pigs in horizontal axis). But with exchange vertical intercept will be 48 and horizontal intercept will be 42.

28 24 12